The specific heat of aluminum is 0.902

J/g

·

o

C

. How much heat is absorbed

by an aluminum

pie tin with a mass of 473 g to raise its temperature

from room temperature (23.00

o

C

) to

oven temperature (375

o

F)?



1
Expert's answer
2021-11-16T12:42:01-0500

"375^oF =(375-32)×(\frac{5}{9})^oC

=190.5^oC"


Increase in temperature "=190.5-23\n\n=167.5"

Required Heat= "0.902\u00d7167.5\u00d7473 J\n\n=71463.205 J"